OpenStudy (anonymous):

The Boardman family walked two over five of a mile in two over seven of an hour. What is their unit rate in miles per hour

OpenStudy (anonymous):

two-sevenths over two-fifths equals one and two-fifths miles per hour two-sevenths over two-fifths equals five over seven mile per hour two-fifths over two-sevenths equals one and two-fifths miles per hour two-fifths over two-sevenths equals five over seven mile per hour

OpenStudy (misty1212):

HI!!! \[\frac{2}{5}\] of a mile in \[\frac{2}{7}\] of an hour

OpenStudy (misty1212):

is that the question?

OpenStudy (anonymous):

yes

OpenStudy (misty1212):

they want you do divide miles by hours \[\frac{2}{5}\div\frac{2}{7}\] you know how to do that?

OpenStudy (anonymous):

no, not really. Sorry.

OpenStudy (misty1212):

as the math teachers say "invert and multiply" which means flip the second on and go straight across \[\frac{2}{5}\div \frac{2}{7}=\frac{2}{5}\times\frac{7}{2}\]

OpenStudy (misty1212):

cancel the twos, gives you \[\frac{7}{5}\]

OpenStudy (anonymous):

true. But thats not one of my answers.

OpenStudy (misty1212):

maybe they want you to write it as a mixed number, or a decimal

OpenStudy (misty1212):

is \[1\tfrac{2}{5}\] a choice?

OpenStudy (anonymous):

yes

OpenStudy (misty1212):

then pick that one!

OpenStudy (anonymous):

So, it would be 2/5 over 2/7 equals 1 and 2/5

OpenStudy (anonymous):

@misty1212